### v1

@@ 2.2 Impact of Dimension Expansion on AI4Science
To analyze the performance of the input expansion method, we compared our results with the benchmark provided by the PINNacle framework [42], as shown in Fig. 4. The improvement was evidenced by the distribution of the loss values, where the expanded input dimensions resulted in a lower overall loss compared to the raw input dimensions. Our findings revealed that approximately 75% of the PDEs showed improved performance with the expanded input dimensions. This improvement was particularly significant for complex and high-dimensional PDEs, demonstrating the effectiveness of input dimension expansion in enhancing the accuracy and robustness of PINNs (the remaining 5 PDEs showed no significant improvement, so they were not included in Fig. 4).
@@ 4.2 Measuring the Degree of Symmetry Breaking
To quantify the degree of symmetry breaking in neural networks, we developed a metric called the “Replica Distance (RD) Metric”. This metric evaluates the Wasserstein distance between network outputs across different weight configurations (i.e., replicas), providing insights into the extent of symmetry breaking. We applied the RD Metric to the CIFAR-10 dataset using five distinct neural network architectures: SimpleCNN, DropoutCNN, BatchNormCNN, FlipEquivarianceCNN, and RotationEquivarianceCNN (180 degrees), analogous to that in Fig. 7. Here, it is important to note that the CIFAR-10 dataset does not exhibit 180-degree rotational equivariance, which makes the rotation constraint less appropriate.
The test accuracies were recorded for each epoch and are shown in Fig. 8 (A), where we see that the FlipEquivarianceCNN achieved the highest test accuracy, followed by DropoutCNN, BatchNormCNN, SimpleCNN, and RotationEquivarianceCNN.
@@
To measure the degree of symmetry breaking, we used a subset of 100 images from the CIFAR-10 dataset, a scenario that is common in network design because we do not want to use the entire dataset during the design and selection phase. Analogous to the ”Parssi” scheme used to determine the replica symmetry of the spin-glass system [25, 26], we trained each model for 200 epochs under 200 different random seeds, and the final weights were cached. Thus, each model has 200 different weight combinations.
@@
For each model, we paired these 200 different weight combinations and calculated a series of Wasserstein distances, then plotted histograms of these distances and applied Gaussian smoothing to obtain smooth curves. The five models correspond to five curves (Fig. 8 (B)). The peak position of the smoothed histogram was used as the metric. A higher peak position indicates a greater degree of symmetry breaking. The FlipEquivarianceCNN exhibited the highest degree of symmetry breaking, followed by RotationEquivarianceCNN, DropoutCNN, BatchNormCNN, and SimpleCNN.
@@
Our proposed metric suggests a possible way to measure the degree of symmetry of neural networks. By analyzing the Wasserstein distance between weight distributions, we can quantify the extent to which different techniques, such as input dimension expansion, dropout, batch normalization, and equivariance constraints, break symmetry in the parameter space. This metric can guide the design of neural network architectures and optimization strategies, helping to improve model performance and generalization.
@@ 5.2 Challenges and Future Directions
While our results are promising, further research is essential to validate and refine the symmetry breaking metric across different datasets and neural network architectures. Extending the analysis to a broader range of datasets, including CIFAR-100, ImageNet, and other domain-specific datasets, will help us understand how the symmetry metric performs across various types of data and tasks, providing a more comprehensive evaluation of its effectiveness. In fact, our experiments have shown that the worse the original network performs, the more significant the improvement brought by input dimension expansion. Therefore, for some tasks where the network design is already very good, the effect of input dimension expansion is not obvious, such as in sentiment analysis and image coloring tasks.
@@ 6.6 Measuring the Degree of Symmetry Breaking
Pairwise Distance Calculation: The pairwise Wasserstein distance between the weight distributions of the trained models was computed to measure the minimum cost of transporting mass in one distribution to match another.
Histogram and Peak Position Calculation: A histogram of the calculated Wasserstein distances was constructed, and Gaussian smoothing was applied to obtain a smooth curve. The peak position of the smoothed histogram was used as the metric. A higher peak position indicates a greater degree of symmetry breaking.
The Wasserstein distance provides a meaningful metric for comparing the distributions of model weights. By examining the peak position of the smoothed histogram, we can quantify the extent to which symmetry is broken in the trained models. A higher peak position suggests that the models have diverged more significantly in their weight distributions, indicating a higher degree of symmetry breaking.

### v2

@@ 2.2 Impact of Dimension Expansion on AI4Science
To analyze the performance of the input expansion method, we compared our results with the benchmark provided by the PINNacle framework [42], as shown in Fig. 4. The improvement was evidenced by the distribution of the loss values, where the expanded input dimensions resulted in a lower overall loss compared to the raw input dimensions. Our findings revealed that approximately 75% of the PDEs showed improved performance with the expanded input dimensions. This improvement was particularly significant for complex and high-dimensional PDEs, demonstrating the effectiveness of input dimension expansion in enhancing the accuracy and robustness of PINNs (the remaining 5 PDEs showed no significant improvement, so they were not included in Fig. 4, which will be explained towards latter Sec. 5.2).
@@ 4.2 Measuring the Degree of Symmetry Breaking
To quantify the degree of symmetry breaking in neural networks, we developed a metric called the “Replica Distance (RD) Metric”. This metric evaluates the Wasserstein distance between different weight configurations of the network (i.e., replicas), providing insights into the extent of symmetry breaking. We applied the RD Metric to the CIFAR-10 dataset using five distinct neural network architectures: SimpleCNN, DropoutCNN, BatchNormCNN, FlipEquivarianceCNN, and RotationEquivarianceCNN (180 degrees), analogous to that in Fig. 7. Here, it is important to note that the CIFAR-10 dataset does not exhibit 180-degree rotational equivariance, which makes the rotation constraint less appropriate.
After training the test accuracies were recorded for each epoch and are shown in Fig. 8 (A), where we see that the FlipEquivarianceCNN achieved the highest test accuracy, followed by DropoutCNN, BatchNormCNN, SimpleCNN, and RotationEquivarianceCNN.
@@
To measure the degree of symmetry breaking, we used a subset of 100 images from the CIFAR-10 dataset, a scenario that is common in network design because we do not want to use the entire dataset during the design and selection phase. Analogous to the “Parassi” scheme used to determine the replica symmetry of the spin-glass system [25, 26], we trained each model for 200 epochs under 200 different random seeds, and the final weights were cached. Thus, each model has 200 different weight combinations.
@@
For each model, we paired these 200 different weight combinations and calculated a series of Wasserstein distances, then plotted histograms of these distances and applied Gaussian smoothing to obtain smooth curves. These smoothed histograms were normalized to form probability density functions. The five models correspond to five curves (Fig. 8 (B)). The mean value of the probability density was used as the metric. A higher average distance indicates a greater degree of symmetry breaking. The FlipEquivarianceCNN exhibited the highest degree of symmetry breaking, followed by RotationEquivarianceCNN, DropoutCNN, BatchNormCNN, and SimpleCNN.
@@
Our proposed metric suggests a way to measure the degree of symmetry of neural networks. By analyzing the Wasserstein distance between weight distributions, we can quantify the extent to which different techniques, such as input dimension expansion, dropout, batch normalization, and equivariance constraints, break symmetry in the parameter space. This metric can guide the design of neural network architectures and optimization strategies, helping to improve model performance and generalization.
@@ 5.2 Challenges and Future Directions
While our results are promising, further research is still essential to further validate and refine the symmetry breaking metric across different datasets and neural network architectures. Extending the analysis to a broader range of datasets, including CIFAR-100, ImageNet, and other domain-specific datasets, will help us more sufficiently and deeply understand how the symmetry metric performs across various types of data and tasks, providing a more comprehensive evaluation of its effectiveness. In fact, our experiments have shown that the worse the original network performs, the more significant the improvement brought by input dimension expansion. Therefore, for some tasks where the network design is already very good, the effect of input dimension expansion is not obvious, or even tends to certainly degenerate the performance due to the good symmetry breaking properties already possessed by the original network architectures, such as in several cases of PDE solving, sentiment analysis and image coloring tasks.
@@ 6.6 Measuring the Degree of Symmetry Breaking
6.7 Measuring the Degree of Symmetry Breaking
Training Different Networks.
6.8 Measuring the Degree of Symmetry Breaking
Training Different Networks.
@@ 6.8 Measuring the Degree of Symmetry Breaking
Dimensionality Reduction: The weights of each trained model were extracted and flattened. UMAP (Uniform Manifold Approximation and Projection) was then used to reduce the dimensionality of these weight vectors to 100 dimensions. UMAP is a non-linear dimensionality reduction technique that is particularly well-suited for preserving the global structure of high-dimensional data in a lower-dimensional space. This step helps in comparing the models by reducing the complexity of the weight space while retaining the essential features.
Pairwise Distance Calculation: The pairwise Wasserstein distance between the reduced feature representations of the trained models was computed to measure the minimum cost of transporting mass in one distribution to match another. This distance metric provides a meaningful way to compare the distributions of model weights.
Histogram and Mean Value Calculation: A histogram of the calculated Wasserstein distances was constructed, and Gaussian smoothing was applied to obtain a smooth curve. These smoothed histograms were normalized to form probability density functions. The mean value of the probability density was used as the metric. A higher mean value indicates a greater degree of symmetry breaking.
The Wasserstein distance provides a meaningful metric for comparing the distributions of model weights. By examining the mean value of the smoothed histogram, we can quantify the extent to which symmetry is broken in the trained models. A higher mean value suggests that the models have diverged more significantly in their weight distributions, indicating a higher degree of symmetry breaking.